OpenStudy (anonymous):

.What is the 24th term of the arithmetic sequence where a1 = 8 and a9 = 56? 134 140 146 152

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm a_9=a_1+d(9-1) }\) \(\large\color{midnightblue}{ \rm a_9=a_1+8d }\) you know the \(\large\color{midnightblue}{ \rm a_1 }\) and the \(\large\color{midnightblue}{ \rm a_9 }\), solve for the difference

OpenStudy (solomonzelman):

After you solve for the difference, use \(\large\color{midnightblue}{ \rm a_{n}=a_1+d(n-1) }\) , or in your case, use \(\large\color{midnightblue}{ \rm a_{25}=a_1+24d }\)

OpenStudy (solomonzelman):

Tell me if you need more help

OpenStudy (anonymous):

Wait, Where did A25 Come from? o-o

OpenStudy (solomonzelman):

My bad

OpenStudy (solomonzelman):

Well, \(\large\color{midnightblue}{ \rm a_{n}=a_1+d(n-1) }\) in your case would be \(\large\color{red}{ \rm a_{24}=a_1+23d }\)

OpenStudy (solomonzelman):

Have you found the difference yet ?

OpenStudy (anonymous):

I Got 6.

OpenStudy (solomonzelman):

yes good !

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm a_{24}=8+23(6) =? }\)

OpenStudy (solomonzelman):

see what I am doing ?

OpenStudy (anonymous):

146?

OpenStudy (solomonzelman):

Yup, great !

OpenStudy (solomonzelman):

How about a sum of the 24 terms ?

OpenStudy (solomonzelman):

or no ?

OpenStudy (anonymous):

Nah, we just had to find the 24th term c: THANKS!

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm S_{n}=\frac{1}{2}(a_1+a_n) \times n }\)

OpenStudy (solomonzelman):

Okay, whatever ... yw !